Given USD/Brazilian Real = 3.486 USD/Australian Dollar = 1.157 USD/Chinese Yuan = 8.336 What is the...

Given

USD/Brazilian Real = 3.486

USD/Australian Dollar = 1.157

USD/Chinese Yuan = 8.336

What is the Australian Dollar/Chinese Yuan cross rate?

Enter your answer rounded off to THREE decimal points.
Answer #1

AUD / CY = (AUD / USD) * (USD / CY)

Given USD / AUD = 1.157

Thus AUD / USD = 1 / 1.157

= 0.8643

CNY / AUD = (CNY / USD) * (USD / AUD)

= 8.336 * 0.8643

= 7.2048

AUD / CNY is 7.2048